Laura is 4 years older than her brother. Two years ago Laura was three times as old as her brother. How old are they now?
I got for my answer 9 Laura and 5 Brother, using (4+b)=3(b-2). I got the brothers age by using 4+b after I put the two equations together and getting b=5 then adding 4 to 5 for Laura and not adding 4 for the brother. But the book says laura is 8 and the brother is 4

Answer by ankor@dixie-net.com(22740)

Let's do it this way and see what happens, sometimes the book is wrong
:
Laura is 4 years older than her brother.
L = b + 4
Two years ago Laura was three times as old as her brother.
L-2 = 3(b-2)
L - 2 = 3b - 6
L = 3b - 6 + 2
L = 3b - 4
replace L with (b+4)
b + 4 = 3b - 4
4 + 4 = 3b - b
8 = 2b
b = 8/2
b = 4 brother's age
:
I see what you did here, can be done your way but,
2 has to be subtracted from both ages
4+b-2 = 3(b-2)
b + 2 = 3b - 6
2 + 6 = 3b - b